# Attention, Distillation, and Tabularization: Towards Practical Neural Network-Based Prefetching

### 23 Dec 2023 v1

@@ Abstract
Attention-based Neural Networks (NN) have demonstrated their effectiveness in accurate memory access prediction, an essential step in data prefetching. However, the substantial computational overheads associated with these models result in high inference latency, limiting their feasibility as practical prefetchers. To close the gap, we propose a new approach based on tabularization that significantly reduces model complexity and inference latency without sacrificing prediction accuracy. Our novel tabularization methodology takes input as a distilled, yet highly accurate attention-based model for memory access prediction and efficiently converts its expensive matrix multiplications into a hierarchy of fast table lookups. As an exemplar of the above approach, we develop DART, a prefetcher comprised of a simple hierarchy of tables. With a modest 0.09 drop in F1-score, DART reduces 99.99% of arithmetic operations from the large attention-based model and 91.83% from the distilled model. DART accelerates the large model inference by 170$\times$ and the distilled model by 9.4$\times$. DART has comparable latency and storage costs as state-of-the-art rule-based prefetcher BO but surpasses it by 6.1% in IPC improvement, resulting in a 37.6% speed-up. DART outperforms state-of-the-art NN-based prefetchers TransFetch by 33.1% and Voyager by 37.2% in terms of IPC improvement, primarily due to its low prefetching latency.
@@ I Introduction
The central challenge in deploying NN-based prefetchers lies in their significant computational demands. Firstly, the quantity of parameters in well-performing NN models impose tremendous system resource requirements, potentially demanding more resources than the application that it was set out to prefetch for. Secondly, these models’ complex nature leads to slow inference, causing untimely data prefetching. Lastly, the sheer volume of arithmetic operations inherent to NNs consumes both substantial resources and energy.
In light of this challenge, motivated by the success of table-based prefetchers, we propose a novel approach that transfers knowledge from a large attention-based neural network to a compact hierarchy of tables for more practical NN-based data prefetching. The key steps of the proposed approach include: 1) training of a large attention-based model for memory access prediction, 2) reducing model complexity via knowledge distillation, and 3) implementing tabularization to transform the distilled model into tables. We develop a prefetcher DART (Distilling Attention-based neuRal network to Tables) as an exemplar of this approach. The predictor in DART is simple and elegant: a hierarchy of tables acquired from the above steps. DART couples the practicality inherent in table-based prefetchers with the precision offered by NN predictors. We summarize our main contributions as follows:
@@
  • We evaluate the prediction performance of DART. With a 0.09 drop of F1-score, DART reduces 99.99% of arithmetic operations from the large attention-based model and 91.83% from the distilled model, accelerating the two models by 170$\times$ and $\times$, respectively.
  • We evaluate the prefetching performance of DART. DART achieves a 37.6% IPC improvement with storage and latency costs comparable to the state-of-the-art rule-based prefetcher BO, surpassing it by 6.1%, and outperforms state-of-the-art NN-based prefetcher TransFetch by 33.1% and Voyager by 37.2%.
@@ VI-A Data Preparation
Segmented address input. We dissects a block address into $S=\lceil\frac{p}{c}\rceil+1$ segments for $p$-bit page address and $c$-bit block index, each housing $c$ bits. This process maps a memory address to a $s$-dimension vector, thereby mapping a $T$-length sequence to a $T\times S$ matrix for attention-based model input.
@@ VI-E1 Layer-Wise Tabularization
17:     else if $\mathcal{M}[i]$ is a LayerNrom then
18:         Push LayerNrom parameters and operations to $\mathcal{T}$
@@ VII-A1 Simulator
Consistent with existing literature, we use ChampSim [46] to generate traces and assess our approach. The simulation parameters are detailed in Table III, with prefetchers simulated at the last-level cache (LLC).
@@ VII-A2 Benchmark
We evaluate DART on widely-used benchmarks: SPEC CPU 2006 [47] and SPEC CPU 2017 [48]. Table IV details LLC trace statistics extracted from ChampSim.
@@ VII-A4 Metrics
We evaluate multi-label memory prediction using F1-score [50] and prefetching with prefetch accuracy, coverage, and IPC improvement [51].
VII-B Evaluation of Multi-Label Knowledge Distillation
@@
VII-C Evaluation of Tabularization
We assess layer-wise tabularization performance under varying table configurations, such as the number of prototypes ($K$) and subspaces ($C$), without fine-tuning. Our model structure remains fixed as the DART model (see Table V), with the only variation being either $K$ or $C$. Figures 7 and 8 illustrate how adjusting $K$ and $C$ impacts DART’s F1-score. The results reveal that increasing $K$ significantly boosts prediction performance when $K$ exceeds 128, with $K=1024$ achieving a 10.9% higher F1 score compared to $K=16$. In contrast, the impact of higher $C$ on F1-score is less pronounced, with $C=8$ outperforming $C=1$ by 6.6%. While higher values of both $K$ and $C$ generally enhance tabularization performance, it’s important to note that this comes at a substantial increase in latency and storage costs. Figure 9 demonstrates a linear relationship between model latency and $\log(K)$ as well as $\log(C)$, while storage costs exhibit exponential growth.
@@ VII-C Evaluation of Tabularization
VII-D Evaluation of Fine-Tuning
@@
VII-E Evaluation of Prefetching
VII-E1 Prefetcher Configurations

### 22 Feb 2024 v3

@@ Abstract
Attention-based Neural Networks (NN) have demonstrated their effectiveness in accurate memory access prediction, an essential step in data prefetching. However, the substantial computational overheads associated with these models result in high inference latency, limiting their feasibility as practical prefetchers. To close the gap, we propose a new approach based on tabularization that significantly reduces model complexity and inference latency without sacrificing prediction accuracy. Our novel tabularization methodology takes input as a distilled, yet highly accurate attention-based model for memory access prediction and efficiently converts its expensive matrix multiplications into a hierarchy of fast table lookups. As an exemplar of the above approach, we develop DART, a prefetcher comprised of a simple hierarchy of tables. With a modest 0.09 drop in F1-score, DART reduces 99.99% of arithmetic operations from the original attention-based model and 91.83% from the distilled model. DART accelerates the large model inference by 170$\times$ and the distilled model by 9.4$\times$. DART has comparable latency and storage costs as state-of-the-art rule-based prefetcher BO but surpasses it by 6.1% in IPC improvement. DART outperforms state-of-the-art NN-based prefetchers TransFetch by 33.1% and Voyager by 37.2% in terms of IPC improvement, primarily due to its low prefetching latency.
@@ I Introduction
The central challenge in deploying NN-based prefetchers lies in their significant computational demands. Firstly, the quantity of parameters in well-performing NN models impose tremendous system resource requirements, potentially demanding more resources than the application that it was set out to prefetch for. Secondly, these models’ complex nature leads to slow inference, causing untimely data prefetching. Lastly, the sheer volume of arithmetic operations inherent to NNs consumes both substantial resources and energy. Though existing works use pruning [20], quantization [21], and hardware parallel implementations [22] to accelerate NN inference, these techniques still require a large number of matrix multiplications during inference.
In light of this challenge, motivated by the success of table-based prefetchers, we propose a novel approach that transfers knowledge from a large attention-based neural network to a compact hierarchy of tables for more practical NN-based data prefetching, eliminating all matrix multiplications in model inference. The key steps of the proposed approach include: 1) training of a large attention-based model for memory access prediction, 2) reducing model complexity via knowledge distillation, and 3) implementing tabularization to transform the distilled model into tables.
In developing the above approach, we tackle several key sub-problems that emerge. 1) How to map layers of neural networks with various operations to table lookups? We design kernels that not only store the precomputed matrix multiplications, but also manage challenges including bias incorporation, table size constraints, and the integration of activation functions between operations. 2) How to reduce the critical path of a model to satisfy a given latency constraint? We quantitatively analyze the table-based model’s inference latency and compress the NN model using a novel knowledge distillation approach for multi-label classification. 3) How to address the error accumulation when mapping multiple NN layers to tables? We propose a novel fine-tuning approach that trains the table to imitate the NN layer output rather than merely approximating dot products between layer inputs and weights. In this way, we mitigate performance degradation due to the approximation as the number of layers increases.
We develop a prefetcher DART (Distilling Attention-based neuRal network to Tables) as an exemplar of this approach. The predictor in DART is simple and elegant: a hierarchy of tables acquired from the above steps. DART couples the practicality inherent in table-based prefetchers with the precision offered by NN predictors. We summarize our main contributions as follows:
@@
  • We evaluate the prediction performance of DART. With a 0.09 drop of F1-score, DART reduces 99.99% of arithmetic operations from the large attention-based model and 91.83% from the distilled model, accelerating the two models by 170$\times$ and 9.4$\times$, respectively.
  • We evaluate the prefetching performance of DART on a variety of workloads from SPEC 2006 and SPEC 2017. DART achieves a 37.6% IPC improvement, outperforming state-of-the-art rule-based prefetcher BO by 6.1% with comparable latency and storage costs. It also surpasses state-of-the-art NN-based prefetchers TransFetch and Voyager by 33.1% and 37.2%, respectively.
@@ VI-A Data Preparation
Segmented address input. We dissect a block address into $S=\lceil\frac{p}{c}\rceil+1$ segments for $p$-bit page address and $c$-bit block index, each housing $c$ bits. This process maps a memory address to an $s$-dimensional vector, transforming a $T$-length sequence to a $T\times S$ matrix for attention-based model input.
@@ VI-E1 Layer-Wise Tabularization
17:     else if $\mathcal{M}[i]$ is a LayerNorm then
18:         Push LayerNorm parameters and operations to $\mathcal{T}$
@@ VII-A1 Simulator
Consistent with existing literature, we use ChampSim [47] to generate traces and assess our approach. ChampSim simulates a modern heterogeneous multi-core system with an arbitrary memory hierarchy. The simulation parameters are detailed in Table III, with prefetchers simulated at the last-level cache (LLC).
@@ VII-A2 Benchmark
We evaluate DART on widely-used benchmarks: SPEC CPU 2006 [48] and SPEC CPU 2017 [49]. Table IV details the workloads we use and the LLC trace statistics extracted using ChampSim. The workloads consist of a diverse number of deltas and pages, which are representative to various scenarios.
@@ VII-A4 Metrics
We evaluate multi-label memory access prediction using F1-score [51]. We evaluate the prefetching performance through simulation on ChampSim. In the simulation runs, we record the cache performance and report the prefetch accuracy, coverage, and IPC improvement [52]. The running time of the workloads on ChampSim cannot indicate the direct application performance. Instead, the IPC performance is reported based on the number of instructions and the cycles of simulation, which accurately indicates the prefetcher performance.
VII-B Memory Access Patterns
Figure 7 shows the diverse memory access patterns of the benchmark applications in three scaled dimensions: instruction ID, page address, and block delta of consecutive accesses. The performance of the original large attention-based memory access prediction model is shown in Table VI as teacher models. The distribution of memory access patterns impacts the model prediction. We observe that applications showing fewer deltas are easier to predict compared with applications with more deltas under a similar number of pages, e.g., 602.gcc vs 605.mcf. We also observe that under a comparable number of deltas, applications with a smaller number of pages are easier to predict, e.g., 410.bwaves vs 433.milc.
Figure 7: Visualization of memory access patterns in the benchmark applications. The patterns show diversity in instructions, deltas, and pages, representing various scenarios.
VII-C Evaluation of Multi-Label Knowledge Distillation
@@
VII-D Evaluation of Tabularization
We assess layer-wise tabularization performance under varying table configurations, such as the number of prototypes ($K$) and subspaces ($C$), without fine-tuning. Our model structure remains fixed as the DART model (see Table V), with the only variation being either $K$ or $C$. Figures 8 and 9 illustrate how adjusting $K$ and $C$ impacts DART’s F1-score. The results reveal that increasing $K$ significantly boosts prediction performance when $K$ exceeds 128, with $K=1024$ achieving a 10.9% higher F1 score compared to $K=16$. In contrast, the impact of higher $C$ on F1-score is less pronounced, with $C=8$ outperforming $C=1$ by 6.6%. While higher values of both $K$ and $C$ generally enhance tabularization performance, it’s important to note that this comes at a substantial increase in latency and storage costs. Figure 10 demonstrates a linear relationship between model latency and $\log(K)$ as well as $\log(C)$, while storage costs exhibit exponential growth.
@@ VII-D Evaluation of Tabularization
VII-E Evaluation of Fine-Tuning
@@
VII-F Evaluation of Prefetching
VII-F1 Prefetcher Configurations
@@ VIII Conclusion
Acknowledgment
This work has been supported by the U.S. National Science Foundation (NSF) under grant CNS-2009057 and SaTC-2104264, as well as the DEVCOM Army Research Lab (ARL) under grant W911NF2220159.
Distribution Statement A: Approved for public release. Distribution is unlimited.